OpenStudy (bittuaryan):

Evaluate: int(dx/x^{2}-4x+5)

OpenStudy (bittuaryan):

I mean:\[\int\limits(dx/x^{2}-4x+5)\]

OpenStudy (anonymous):

you can also write this like \[\int\limits (1/x^2-4x+5)dx\]

OpenStudy (anonymous):

\[\int\limits(\frac{ 1 }{ x^2 }-4x+5)dx=-\frac{ 1 }{ x }-2x^2+5x\]

OpenStudy (anonymous):

sorry + C cant forget the constant 'C' since this is an indefinite integral

OpenStudy (bittuaryan):

please clear step 2: (-1/x-2x^2+5x)

OpenStudy (anonymous):

did this help @bittuaryan ?

OpenStudy (bittuaryan):

@sjerman1 How (1/(x^2−4x+5)) is change to ((1/x^2)−4x+5).

OpenStudy (anonymous):

haha well you need to specify with parentheses because i thought only the first term was 1/x^2 ...

OpenStudy (anonymous):

well then in this case since x^2-4x+5 can not be factored we can complete the square to get something that can be used

OpenStudy (anonymous):

atan(-2+x) is the answer

OpenStudy (anonymous):

after completing the square we get 1/(x-2)^2+1

OpenStudy (anonymous):

use u-sub for x-2 and get 1/(u^2+1)

OpenStudy (kropot72):

\[x ^{2}-4x+5=(x-5)(x+1)\]

OpenStudy (anonymous):

sjerman1 is correct

OpenStudy (anonymous):

we know that 1/u^2+1 is arctan so arctan(u)

OpenStudy (anonymous):

then just plug in u=x-2

OpenStudy (anonymous):

atan(x-2)

OpenStudy (anonymous):

I verified in matlab. good job.

OpenStudy (anonymous):

and @fmg78360 has the answer -> arctan(x-2)

OpenStudy (anonymous):

I cheated you did the actual work

OpenStudy (anonymous):

we all did good here. and now we know how to do it!

OpenStudy (abb0t):

did you mean: \[\int\limits \frac{ dx }{ x^2-4x+5 }\]

OpenStudy (anonymous):

yes, @abb0t he did, but i didn't catch that in the beginning.

OpenStudy (abb0t):

Use partial fraction decomposition to solve the integral: \[\int\limits\limits \frac{ 1 }{ x^2-4x+5 }dx \neq \int\limits\limits \frac{ 1 }{ (x-5)(x+1) }\]

OpenStudy (anonymous):

or you can use completing the square like i did

OpenStudy (anonymous):

The integral was solved using the arctan integral identity.

OpenStudy (abb0t):

I didn't see. But complete the square works.